# Seele: A Unified Acceleration Framework for Real-Time Gaussian Splatting

### v2

@@ 5.1 Algorithm
The goal of our contribution-aware rasterization is to pay “computational” attention to significant Gaussians that contribute more to the final pixel values while reducing computations for less important Gaussians. Our observation is that the color values of most pixels are determined by a small fraction of Gaussians, as illustrated in Fig. 4. We characterize the MipNeRF360 dataset [13] and draw the accumulative transparency of each pixel by sorting the Gaussian transparency contribution in descending order. The result in Fig. 4 shows that 1.5% of the top Gaussians contribute to 99% of the final pixel. The remaining Gaussians contribute minimally, primarily adding low-frequency textures. This is not an artifact but an inherent feature of Gaussian representation. Because low values are typically indicative of low-frequency information, whereas high values are often associated with high-frequency details, as shown by the red crosses in Fig. 4.
Idea.
To leverage this observation, we propose a contribution-aware rasterization algorithm, as outlined in Algo. 1. Instead of uniformly assigning the same computation to all Gaussians, our algorithm organizes every pixels into small groups, P. During each iteration, only the leader pixel within a group, which is the centroid of this group, computes the value for a given Gaussian, . If this value falls below a predefined threshold (, to avoid numerical instability), the algorithm skips the color blending of this Gaussian on the remaining pixels within the group, P. This selective skipping largely reduces the workload for less important Gaussians.
@@
Not only does our algorithm reduce the overall computation, but it also inherently achieves higher parallelism by alleviating warp divergence. In the GPU implementation, pixels within a group are assigned to one warp. All pixels within a warp execute in “lockstep” as shown in Fig. 7. For instance, when blending Gaussian , only thread accumulates the insignificant Gaussian into its pixel, while all other threads are masked by GPU and wait for thread to complete.
@@
In contrast, our algorithm only lets thread detect if Gaussian is insignificant. In this case, all threads in this warp skip the color blending of Gaussian . In the case of Gaussian , because it is considered significant, all threads would blend Gaussian instead. This ensures that these pixels either skip or process the same Gaussians collectively. This way, we eliminate branching within a warp.
5.2 Memory Optimization
In our contribution-aware rasterization, we further introduce a fine-grained pipelining that overlaps computation with data fetching from GPU memory to shared memory.
Double Buffering.
Fig. 5 illustrates a simplified GPU design with our double buffering strategy. Specifically, it shows a streaming multiprocessor (SM), in Nvidia’s terminology, which consists of multiple compute cores. Each SM includes several GPU cores and a shared memory, a relatively small on-chip buffer (32-64 KB) for faster data access. Compared to accessing GPU memory, reading data from shared memory is often much faster. Note that, we intentionally ignore other buffers, e.g., L2 caches, in the GPU memory hierarchy to keep the illustration simple.
Our double buffering strategy splits the shared memory into two buffers: a working buffer and a filling buffer. The filling buffer preloads data for the next batch of computations, while the working buffer stores the data required for the current batch. In Fig. 5, the red arrows show that different GPU cores access data from the working buffer for their own computations. Meanwhile, as the green arrow shows, the filling buffer fetches data from GPU memory, preparing for the next batch of computations.
Fine-Grained Pipelining.
With our double buffering strategy, the rasterization pipeline achieves a more fine-grained pipelining compared to the conventional rasterization pipeline in canonical Gaussian splatting. Fig. 6 provides a simplified example of rendering frames within a single warp. In the original algorithm, data loading and computing (rasterization) have to execute sequentially due to the warp-level synchronization. The on-chip data need to be synchronized to ensure the rendering’s correctness. However, the synchronization after compute batch unnecessary since there is no data dependency between computing the current batch and loading data for the next batch. This unnecessary synchronization is because the entire shared memory is allocated to a single batch of computation, forcing synchronization after each batch.
Thus, our pipeline partitions shared memory, allocating only half for each batch of computation. This allows data to be loaded into the second half while the current batch is being processed, as illustrated in Fig. 6.
Specifically, each warp loads the attributes of two Gaussians, with each thread within a warp loading one byte of the Gaussian attributes. A single GPU block, which consists of 8 warps, effectively loads 16 Gaussian points into shared memory. Here, we intentionally keep the batch size small to achieve finer-grained pipelining. For more details on the implementation, please refer to our code, which will be made available upon acceptance of this work.
6 Integrated Fine-Tuning
@@
Note that, both HP and CR modify the rendering pipeline. To regain the rendering quality, our algorithm only requires minimal fine-tuning to achieve better rendering quality. Since both hybrid preprocessing and contribution-aware rasterization do not participate in gradient descent, these two techniques can be integrated into the current training process without any notifications on back-propagation.
Once the 3DGS model is trained by the canonical training process, we convert the model into our view-dependent scene representation as described in Sec. 4. Next, we fix the shared Gaussians and fine-tune only the exclusive Gaussians in each cluster individually for 1000 steps. Since our algorithm always takes clusters (1 nearest cluster and neighboring clusters) during rendering, it naturally accommodates the view discrepancies when shifting between two cluster boundaries. Lastly, we fix the exclusive Gaussians and fine-tune the shared ones for 1000 steps.
7 Evaluation
![[Uncaptioned image]](extracted/6266021/figs/gold_medal.png)
![[Uncaptioned image]](extracted/6266021/figs/silver_medal.png)
CompactGS [32] | 26.99 | 0.797 | 0.247 | 24.18 | 1824.57 | 122.7 | 23.42 | 0.831 | 0.202 | 46.53 | 920.04 | 93.0 | 29.84 | 0.902 | 0.257 | 32.87 | 1294.45 | 104.0 |
EAGLES [20] | 27.14 | 0.805 | 0.239 | 22.30 | 1988.56 | 68.1 | 23.34 | 0.836 | 0.200 | 60.27 | 713.29 | 34.0 | 29.95 | 0.907 | 0.249 | 25.24 | 1555.53 | 61.5 |
AdR-Gaussian [44] | 26.90 | 0.783 | 0.272 | 65.95 | 885.27 | 274.3 | 23.53 | 0.832 | 0.205 | 114.18 | 442.99 | 192.0 | 29.65 | 0.901 | 0.254 | 78.87 | 485.00 | 335.5 |
3DGS [27] | 27.46 | 0.812 | 0.223 | 20.79 | 2168.68 | 710.6 | 23.75 | 0.848 | 0.176 | 41.97 | 1034.37 | 430.9 | 29.59 | 0.902 | 0.244 | 23.94 | 1645.11 | 639.9 |
@@ 7 Evaluation
Seele + LightGaussian | 27.56 | 0.810 | 0.229 |
0.846 | 0.181 |
29.73 | 0.900 | 0.249 |
@@ 7.1 Experimental Setup
In addition to the two proposed techniques in Sec. 3, we also implement the following optimizations. First, we apply fine-grained data fetching when loading Gaussian attributes to better utilize shared memory on GPU. Second, we leverage thread-level parallelism to overlap the execution of fetching Gaussian attributes and computation of alpha blending. Lastly, we replace the compute-intensive exp() function with CUDA intrinsic function __exp(), which generates fewer machine instructions once complied. Note that, all optimizations above are nothing more than engineering hacks to improve the program’s efficiency; we do not claim them as our major contributions.
Dataset and Metrics.
@@
The primary hardware we used is a mobile Ampere GPU on Nvidia AGX Orin SoC, which is the flagship development board in AR/VR and autonomous driving. Meanwhile, we also report the performance numbers on Nvidia Orin NX [5], a low-power embedded computing board, and Nvidia A6000 [3], a powerful workstation GPU.
@@ 7.2 Performance and Quality
Tbl. 1 shows the overall evaluation of Seele on three widely-used 3DGS pipelines: 3DGS [27], MiniSplatting [16], and LightGaussian [15]. Overall, Seele generally achieves better rendering quality on all three datasets across all three evaluation metrics. For instance, on average, Seele improves PSNR and SSIM by 0.28 dB and 0.004, respectively. Fig. 10 and Fig. 11 show the examples of the qualitative comparison between Seele and its corresponding baselines. Visually, there is no quality difference between Seele and the corresponding baselines.
Metrics | Method | Mip-NeRF360 | Tanks&Temples | Deep Blending | ||||||||||
Bicycle | Bonsai | Counter | Flowers | Garden | Kitchen | Room | Stump | Treehill | Train | Truck | Playroom | Drjohnson | ||
PSNR | 3DGS [27] | 25.11 | 32.20 | 29.08 | 21.44 | 27.38 | 31.24 | 31.56 | 26.70 | 22.46 | 22.06 | 25.44 | 29.92 | 29.26 |
Seele + 3DGS | 25.02 | 33.15 | 29.60 | 21.50 | 27.45 | 31.67 | 32.04 | 26.57 | 22.49 | 22.39 | 25.65 | 30.11 | 29.48 | |
MiniSplatting [16] | 24.86 | 31.86 | 28.74 | 21.16 | 26.18 | 31.38 | 31.47 | 26.78 | 22.63 | 21.37 | 25.00 | 30.42 | 29.66 | |
Seele + MiniSplatting | 25.21 | 32.86 | 29.31 | 21.38 | 26.96 | 32.11 | 31.69 | 26.97 | 22.79 | 22.02 | 25.46 | 30.37 | 29.67 | |
LightGaussian [15] | 25.13 | 32.20 | 28.98 | 21.41 | 27.01 | 31.14 | 31.87 | 26.66 | 22.53 | 22.16 | 25.47 | 30.03 | 29.44 | |
Seele + LightGaussian | 24.94 | 32.72 | 29.32 | 21.44 | 27.06 | 31.43 | 32.03 | 26.52 | 22.54 | 22.30 | 25.52 | 30.03 | 29.43 | |
SSIM | 3DGS [27] | 0.745 | 0.945 | 0.915 | 0.589 | 0.857 | 0.931 | 0.927 | 0.768 | 0.634 | 0.817 | 0.879 | 0.902 | 0.902 |
Seele + 3DGS | 0.745 | 0.950 | 0.919 | 0.593 | 0.858 | 0.934 | 0.929 | 0.765 | 0.634 | 0.821 | 0.882 | 0.903 | 0.903 | |
MiniSplatting [16] | 0.750 | 0.948 | 0.913 | 0.592 | 0.827 | 0.932 | 0.930 | 0.787 | 0.649 | 0.803 | 0.872 | 0.908 | 0.908 | |
Seele + MiniSplatting | 0.763 | 0.953 | 0.920 | 0.607 | 0.842 | 0.937 | 0.932 | 0.795 | 0.655 | 0.815 | 0.878 | 0.908 | 0.908 | |
LightGaussian [15] | 0.743 | 0.941 | 0.907 | 0.585 | 0.840 | 0.926 | 0.925 | 0.769 | 0.630 | 0.806 | 0.878 | 0.902 | 0.899 | |
Seele + LightGaussian | 0.743 | 0.945 | 0.911 | 0.590 | 0.843 | 0.929 | 0.926 | 0.767 | 0.633 | 0.811 | 0.880 | 0.902 | 0.898 | |
LPIPS | 3DGS [27] | 0.245 | 0.181 | 0.183 | 0.359 | 0.124 | 0.118 | 0.197 | 0.245 | 0.347 | 0.203 | 0.148 | 0.247 | 0.241 |
Seele + 3DGS | 0.238 | 0.176 | 0.177 | 0.353 | 0.118 | 0.113 | 0.193 | 0.241 | 0.343 | 0.194 | 0.141 | 0.243 | 0.237 | |
MiniSplatting [16] | 0.256 | 0.166 | 0.176 | 0.353 | 0.171 | 0.115 | 0.184 | 0.237 | 0.344 | 0.223 | 0.150 | 0.245 | 0.240 | |
Seele + MiniSplatting | 0.240 | 0.161 | 0.168 | 0.340 | 0.155 | 0.109 | 0.181 | 0.223 | 0.332 | 0.214 | 0.144 | 0.245 | 0.239 | |
LightGaussian [15] | 0.255 | 0.192 | 0.201 | 0.372 | 0.148 | 0.130 | 0.204 | 0.247 | 0.367 | 0.227 | 0.150 | 0.251 | 0.249 | |
Seele + LightGaussian | 0.248 | 0.186 | 0.194 | 0.366 | 0.142 | 0.125 | 0.200 | 0.243 | 0.358 | 0.219 | 0.143 | 0.250 | 0.248 | |
FPS on Orin | 3DGS [27] | 12.18 | 33.93 | 24.61 | 27.84 | 16.57 | 19.65 | 22.15 | 23.00 | 22.28 | 43.67 | 40.39 | 28.94 | 20.41 |
Seele + 3DGS | 40.15 | 98.47 | 67.76 | 71.91 | 46.83 | 54.63 | 86.06 | 45.00 | 68.86 | 139.28 | 118.08 | 93.38 | 80.70 | |
MiniSplatting [16] | 93.41 | 59.33 | 45.65 | 96.50 | 91.85 | 50.32 | 67.71 | 95.34 | 96.43 | 149.66 | 137.40 | 133.58 | 108.93 | |
Seele + MiniSplatting | 176.80 | 116.21 | 89.38 | 177.59 | 162.34 | 88.62 | 121.36 | 161.34 | 184.57 | 264.27 | 271.89 | 233.54 | 175.84 | |
LightGaussian [15] | 18.71 | 43.64 | 35.29 | 41.92 | 28.43 | 29.11 | 29.01 | 36.43 | 31.90 | 74.57 | 58.57 | 51.89 | 36.96 | |
Seele + LightGaussian | 61.43 | 102.4 | 81.44 | 88.64 | 62.33 | 69.09 | 75.58 | 80.80 | 82.33 | 209.73 | 163.67 | 143.35 | 120.88 | |
FPS on Orin NX | 3DGS [27] | 1.74 | 5.03 | 3.46 | 4.35 | 2.53 | 2.73 | 3.23 | 3.59 | 3.33 | 6.34 | 5.75 | 4.50 | 3.06 |
Seele + 3DGS | 6.62 | 14.68 | 10.81 | 12.04 | 7.02 | 7.71 | 12.07 | 9.44 | 10.00 | 21.93 | 17.76 | 16.59 | 13.40 | |
MiniSplatting [16] | 9.32 | 8.05 | 6.16 | 10.01 | 10.26 | 6.82 | 8.88 | 10.36 | 10.32 | 21.68 | 19.57 | 18.76 | 15.28 | |
Seele + MiniSplatting | 20.25 | 16.38 | 13.02 | 20.73 | 20.05 | 13.17 | 18.39 | 19.48 | 20.94 | 43.39 | 40.29 | 36.05 | 29.42 | |
LightGaussian [15] | 3.25 | 6.60 | 4.71 | 5.71 | 3.93 | 4.30 | 3.58 | 4.91 | 4.31 | 10.28 | 8.04 | 7.00 | 4.92 | |
Seele + LightGaussian | 10.14 | 18.03 | 13.11 | 15.85 | 10.33 | 11.12 | 13.91 | 12.46 | 13.27 | 31.83 | 24.13 | 22.45 | 18.67 | |
FPS on A6000 | 3DGS [27] | 78.47 | 212.49 | 147.78 | 186.60 | 107.79 | 123.18 | 142.96 | 160.67 | 148.65 | 190.77 | 204.83 | 193.24 | 130.62 |
Seele + 3DGS | 217.39 | 499.75 | 401.61 | 398.57 | 233.92 | 289.02 | 449.64 | 336.36 | 353.98 | 573.72 | 504.29 | 556.79 | 480.08 | |
MiniSplatting [16] | 518.67 | 348.55 | 274.80 | 532.77 | 508.65 | 304.32 | 407.17 | 546.45 | 529.94 | 685.40 | 324.36 | 768.64 | 645.16 | |
Seele + MiniSplatting | 809.72 | 558.97 | 464.90 | 798.08 | 720.46 | 478.47 | 640.61 | 725.16 | 824.40 | 1081.08 | 1137.66 | 1124.86 | 952.38 | |
LightGaussian [15] | 129.25 | 260.08 | 193.09 | 234.69 | 165.84 | 179.21 | 190.95 | 217.20 | 168.69 | 343.64 | 305.53 | 285.31 | 198.85 | |
Seele + LightGaussian | 334.78 | 594.53 | 449.03 | 527.70 | 342.35 | 401.61 | 487.57 | 424.27 | 449.64 | 827.81 | 694.44 | 741.29 | 589.28 | |
#Inst.() | 3DGS [27] | 3333.20 | 1390.33 | 1948.76 | 1456.05 | 2607.93 | 2446.63 | 2668.26 | 1806.44 | 1860.57 | 944.96 | 1123.78 | 1212.77 | 2077.46 |
Seele + 3DGS | 1101.41 | 523.39 | 613.71 | 594.67 | 1017.66 | 990.87 | 746.22 | 732.72 | 683.65 | 320.99 | 391.32 | 356.96 | 486.52 | |
MiniSplatting [16] | 727.18 | 831.48 | 913.19 | 652.21 | 700.11 | 1032.53 | 962.47 | 679.40 | 683.06 | 312.32 | 347.06 | 347.31 | 418.42 | |
Seele + MiniSplatting | 399.97 | 459.61 | 482.96 | 359.97 | 392.56 | 565.16 | 492.82 | 403.60 | 371.00 | 169.88 | 174.92 | 205.22 | 243.94 | |
LightGaussian [15] | 2008.94 | 1060.64 | 1291.92 | 1142.53 | 1782.13 | 1630.33 | 1928.39 | 1467.85 | 1488.80 | 582.87 | 816.55 | 851.42 | 1344.23 | |
Seele + LightGaussian | 731.61 | 411.15 | 504.45 | 458.44 | 740.96 | 693.52 | 625.09 | 601.90 | 541.49 | 219.18 | 300.37 | 276.82 | 371.69 | |
Mem. (MB) | 3DGS [27] | 1320.0 | 295.7 | 276.2 | 772.4 | 1305.1 | 410.5 | 352.0 | 942.9 | 720.8 | 257.5 | 604.3 | 543.3 | 736.6 |
Seele + 3DGS | 706.9 | 145.4 | 166.6 | 382.3 | 728.3 | 295.8 | 204.3 | 433.9 | 364.2 | 135.5 | 279.6 | 387.0 | 414.3 | |
MiniSplatting [16] | 157.6 | 147.0 | 130.7 | 142.4 | 176.0 | 142.3 | 121.5 | 157.8 | 136.7 | 85.0 | 81.5 | 117.4 | 192.3 | |
Seele + MiniSplatting | 107.8 | 91.7 | 106.9 | 95.7 | 126.1 | 122.4 | 98.0 | 114.8 | 95.4 | 62.6 | 57.7 | 104.9 | 154.3 | |
LightGaussian [15] | 106.7 | 24.1 | 23.8 | 64.2 | 94.4 | 34.6 | 28.3 | 89.4 | 68.7 | 23.5 | 44.2 | 38.9 | 66.0 | |
Seele + LightGaussian | 65.3 | 15.0 | 19.4 | 37.2 | 66.3 | 30.6 | 21.9 | 55.3 | 41.2 | 14.7 | 26.6 | 33.6 | 49.2 |
@@
Seele consistently achieves speedups across all algorithms. On average, Seele achieves 3.2, 1.8, and 2.7 speedup on 3DGS, MiniSplatting and LightGaussian, respectively. The higher speedup observed in 3DGS compared to the other two algorithms can be attributed to the denser nature of 3DGS. This density allows Seele to better separate irrelevant Gaussians contributed to different views using hybrid preprocessing introduced in Sec. 4, resulting in greater speedup. Despite that, sparse models like MiniSplatting still benefit from hybrid preprocessing due to the view-dependent redundancy, i.e., different views require different Gaussians to render. Tbl. 1 also shows the average number of executed instructions of each algorithm using NVIDIA Nsight Compute [4], Seele indeed reduces the overall computation. Tbl. 3 further dissects the contributions.
Tbl. 2 shows the speedup of Seele on additional GPUs, Nvidia Orin NX and Nvidia A6000. The results show that the optimizations proposed in Seele are not tied to a specific GPU architecture, exhibiting good generality across different hardware platforms. However, we observe that our optimization achieves a higher speedup on lower-power GPUs, as shown in Tbl. 2. For example, Seele +3DGS achieves a speedup on the Orin NX, while it achieves a speedup on the A6000. Primarily, it is because low-end GPUs often have restricted hardware resources, i.e., compute units, GPU bandwidth, etc.
Seele not only speeds up the overall rendering process but also reduces overall GPU memory consumption. Here, we exclusively focus on the GPU memory contributed by the Gaussian points, i.e., runtime model weights. Overall, results show that Seele achieves 32.3% reduction in runtime model size across three different datasets. Even for sparse models like MiniSplatting, Seele still saves 23.2% of model size compared to the baselines.

### v3

@@ 3.4 Contribution-Aware Rasterization, CR
The goal of our contribution-aware rasterization is to pay “computational” attention to significant Gaussians that contribute more to the final pixel values while reducing computations for less important Gaussians. Our observation is that the color values of most pixels are determined by a small fraction of Gaussians, as illustrated in Fig. 4. We characterize the MipNeRF360 dataset [barron2021mip] and draw the cumulative transparency of each pixel by sorting the Gaussian transparency contribution in descending order. The result in Fig. 4 shows that 1.5% of the top Gaussians contribute to 99% of the final pixel, while remaining Gaussians contribute minimally due to the lower sampling . Furthermore, these low contribution samples typically occur in Gaussians with larger variances or at locations farther from the mean, corresponding to the low-frequency components and sharing low gradients and good spatial locality. More rigorous analysis of the correlation between Gaussians and frequency is shown in the supplementary.
Algorithm.
To leverage the observation, we propose a contribution-aware rasterization algorithm to identify insignificant samples and rearrange computational resources, as outlined in Algo. 1. Instead of uniformly assigning the same computation to all Gaussians, our algorithm organizes every pixels into a small group, P. During each iteration, only the leader pixel within a group, which is the centroid of this group, computes the value for a given Gaussian, . If this value falls below a predefined threshold (, to avoid numerical instability), the algorithm skips the color blending of this Gaussian on all pixels within P. We set to 2 to balance rendering efficiency and quality, and analyze the upper bound on the blending error introduced by this skipping in the appendix.
@@
Not only does our algorithm reduce the overall computation, but it also inherently achieves higher parallelism by alleviating warp divergence. In the GPU implementation, pixels within a group are assigned to one warp. All pixels within a warp execute in “lockstep” as shown in Fig. 5. For instance, when blending Gaussian , only thread accumulates the insignificant Gaussian into its pixel, while other threads wait for thread to complete.
@@
In contrast, our algorithm only lets thread detect if Gaussian is insignificant. In this case, all threads in this warp skip the color blending of Gaussian . In the case of Gaussian , if it is considered significant, all threads would blend Gaussian . This ensures that these pixels either skip or process the same Gaussians collectively. This way, we eliminate divergence within a warp.
3.5 Integrated Fine-Tuning
@@
Note that, both HP and CR modify the rendering pipeline and might impact the view consistency. To maintain a better view consistency across frames, our algorithm only requires minimal fine-tuning to achieve better rendering quality. Since both hybrid preprocessing and contribution-aware rasterization do not participate in gradient descent, these two techniques can be integrated into the current training process without any modifications to back-propagation.
Once the 3DGS model is trained by the canonical training process, we convert the model into our view-dependent scene representation in Sec. 3.3. Next, we fine-tune the model for additional 1000 steps with our loss function,
where is the original loss function in 3DGS and is the view consistency loss. Here, we use F LIP score [andersson2020flip] with 7 frames as our metric. is set to be 0.1.
and
denote the best and second-best results among all methods, respectively.| Dataset | Mip-NeRF360 | Tanks&Temples | Deep Blending | ||||||||||||||||||||||||||
| Metrics | Quality | Efficiency | Quality | Efficiency | Quality | Efficiency | |||||||||||||||||||||||
| PSNR | SSIM | LPIPS | FPS |
PSNR | SSIM | LPIPS | FPS |
PSNR | SSIM | LPIPS | FPS |
| 3DGS [kerbl20233d] | 27.46 | 0.812 | 0.223 | 20.79 | 2168.68 | 710.6 | 23.75 |
0.848 |
0.176 | 41.97 | 1034.37 | 430.9 | 29.59 | 0.902 | 0.244 | 23.94 | 1645.11 | 639.9 | |||||||||||
@@
52.4 |
| Seele + LightGaussian | 27.56 | 0.810 | 0.229 | 76.36 | 589.85 |
39.1
23.91
0.846 | 0.181 | 183.86 | 259.78 |
20.7
29.73 | 0.900 | 0.249 | 131.16 | 324.25 |
41.4
| AdR-Gaussian [wang2024adr] | 26.92 | 0.786 | 0.267 | 65.95 | 885.27 | 288.6 | 23.42 | 0.835 | 0.202 | 114.18 | 442.99 | 203.0 | 29.68 | 0.901 | 0.255 | 78.87 | 485.00 | 330.5 | |||||||||||
| Seele + AdR-Gaussian | 27.19 | 0.790 | 0.261 |
96.25
327.26
147.6 | 23.78 | 0.838 | 0.196 |
193.83
154.87
106.2 | 29.64 | 0.899 | 0.254 |
141.87
178.16
127.7 | |||||||||||
4 Evaluation
@@ 4.1 Experimental Setup
In addition to the two proposed techniques in Sec. 3, we also implement the following optimizations. First, we apply fine-grained data fetching when loading Gaussian attributes to better utilize shared memory on GPU. Second, we leverage thread-level parallelism to overlap the execution of fetching Gaussian attributes and computation of alpha blending. Lastly, we replace the compute-intensive exp() function with CUDA intrinsic function __exp(), which generates fewer machine instructions once compiled. Note that, all optimizations above are nothing more than engineering hacks to improve the program’s efficiency; we do not claim them as our major contributions.
Datasets.
To show the efficiency and robustness of Seele, we evaluate on three datasets: Mip-NeRF360 [barron2021mip], Tank&Temple [knapitsch2017], and DeepBlending [hedman2018deep]. For quality evaluation, we use PSNR, SSIM, and LPIPS. We also report FPS, the number of executed instructions (#Inst.), and the runtime model size as our performance metrics.
@@
To show the general applicability of our optimization technique, we apply our framework to four state-of-the-art 3DGS algorithms: 3DGS [kerbl20233d], MiniSplatting [fang2024mini], LightGaussian [fan2023lightgaussian], and AdR-Gaussian [wang2024adr].
@@
The primary hardware we used is a mobile Ampere GPU on Nvidia AGX Orin SoC, which is the flagship development board in AR/VR and autonomous driving. Meanwhile, we also report the performance numbers on Nvidia Orin NX, a low-power embedded computing board, and Nvidia A6000, a powerful workstation GPU.
@@ 4.2 Performance and Quality
Tbl. 1 shows the overall evaluation of Seele on four widely-used 3DGS pipelines: 3DGS [kerbl20233d], MiniSplatting [fang2024mini], LightGaussian [fan2023lightgaussian], and AdR-Gaussian [wang2024adr]. Generally, Seele achieves better rendering quality on all three datasets across all three evaluation metrics. For instance, on average, Seele improves PSNR and SSIM by 0.28 dB and 0.004, respectively. More qualitative comparisons are shown in the supplementary.
We also measure the view consistency metrics, F LIP1 and F LIP7 [andersson2020flip] in Tbl. 2. With the loss function proposed in Sec. 3.5, Seele achieves better view consistency against the corresponding baselines.
| Dataset | Mip-NeRF360 | Tanks&Temples | Deep Blending | |||
| GPU | Orin-NX | A6000 | Orin-NX | A6000 | Orin-NX | A6000 |
| 3DGS | 3.06 | 134.45 | 6.03 | 197.55 | 3.64 | 155.87 |
| Seele + 3DGS | 9.42 | 328.98 | 19.63 | 536.77 | 14.82 | 515.60 |
| MiniSplatting | 9.62 | 414.02 | 20.57 | 440.34 | 16.84 | 701.51 |
| Seele + MiniSplatting | 17.48 | 640.16 | 41.78 | 1108.65 | 32.40 | 1031.46 |
| LightGaussian | 4.39 | 185.95 | 9.02 | 323.47 | 5.78 | 234.36 |
| Seele + LightGaussian | 12.72 | 431.99 | 27.45 | 755.29 | 20.38 | 656.60 |
| AdR-Gaussian | 10.30 | 383.35 | 16.65 | 456.12 | 13.43 | 478.45 |
| Seele + AdR-Gaussian | 16.70 | 524.88 | 31.84 | 762.71 | 19.57 | 762.45 |
@@
With comparable rendering quality, Seele consistently achieves speedups across all algorithms. On average, Seele achieves 3.2, 1.8, 2.7, and 1.7 speedup on 3DGS, MiniSplatting, LightGaussian, and AdR-Gaussian respectively. The higher speedup observed in 3DGS compared to the other algorithms can be attributed to the denser nature of 3DGS. This density allows Seele to better separate irrelevant Gaussians contributed to different views using hybrid preprocessing introduced in Sec. 3.3, resulting in greater speedup. Despite that, sparse models like MiniSplatting still benefit from hybrid preprocessing due to the view-dependent redundancy, i.e., different views require different Gaussians to render. Tbl. 1 also shows the average number of executed instructions of each algorithm using NVIDIA Nsight Compute, Seele indeed reduces the executed instructions, thus overall computation. Tbl. 4 further dissects the contributions of our techniques.
Tbl. 3 shows the speedup of Seele on additional GPUs, Nvidia Orin NX and Nvidia A6000. The results show that the optimizations proposed in Seele are not tied to a specific GPU architecture, exhibiting good generality across different hardware platforms. However, we observe that our optimization achieves a higher speedup on lower-power GPUs, as shown in Tbl. 3. Primarily, low-end GPUs often have more restricted hardware resources.
Seele not only speeds up the overall rendering process but also reduces overall GPU memory consumption. Here, we exclusively focus on the GPU memory contributed by the Gaussian points, i.e., runtime model weights. Overall, results show that Seele achieves 39.1% reduction in runtime model size across three different datasets. Even for sparse models like MiniSplatting, Seele still saves 23.2% of model size compared to the baselines.